# Multi-site gates for state preparation in quantum simulation of the Bose Hubbard Model

### v1

@@ Abstract
State preparation is one of the most important challenges in analog quantum simulation. The most common approach, adiabatic state preparation, performs poorly in gapless systems and cannot efficiently cross phase boundaries [1, 2, 3]. Driven dissipative approaches are promising, but require specialized hardware [4, 5]. Here we present a coherent approach to producing the ground state of the Bose Hubbard model, based upon simple multi-site gates. These are more flexible than the two-site gates used in digital quantum simulation, and allow us to produce the exact ground state, without any Trotter error [6, 7, 8, 9]. Related ideas have been proposed for producing states with applications to quantum information theory [10, 11, 12]. Although our approach can be used in a broad range of architectures [6] it is particularly appealing for simulation based on transmon arrays, as all of the necessary resources are found in the current generation of NISQ era quantum computers[13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Here $d_{i},d_{i}^{\dagger}$ are ladder operators which change the number of excitations present in the transmon on site $i$, $g_{ij}$ parameterizes the strength of the coupling between sites $i$ and $j$, $\delta_{i}$ is the detuning, and $\eta$ encodes the transmon nonlinearity. In modern hardware both $g_{ij}$ and $\delta_{i}$ are tunable, and can be made functions of time [21]. If we interpret $d_{i}$ as the annihilation operator for a boson, Eq. (1) is the Bose Hubbard model. In that interpretation, $g_{ij}$ is a hopping matrix element, $\delta_{i}$ an on-site potential, and $\eta$ an on-site interaction – more commonly written as $t_{ij},V_{i},$ and $U$. The Bose Hubbard model is one of the iconic models of strongly interacting quantum matter, displaying fascinating emergent properties, including a phase transition between a superfluid and insulating state. While there are efficient classical methods to calculate any static properties of this model [25], its dynamics are hard to calculate on a classical computer. Thus this is a natural setting to demonstrate quantum advantage [26].
@@ II.1 Single Particle Physics
We begin by presenting a one dimensional model, where $g_{ij}$ vanishes unless $j=i+1$. We will choose these matrix elements so that the hopping part of the Hamiltonian can be mapped onto a SU(2) spin matrix. The structure of spin matrices allow us to construct a gate which produces the superfluid ground state of the Bose Hubbard model.
@@ II.2 Many-Body Physics of our 1D model
Figure 2: Fidelity $|\langle\psi_{X}|\psi\rangle|^{2}$ of state preparation using the pulse sequence in Sec. II.2 for the 1D Bose Hubbard model. Our target state, $\psi_{X}$, is the ground state of the Bose Hubbard Hamiltonian $H_{X}$ in Eq. (10, while $|\psi\rangle=U_{Z}(\varphi)U_{X}(\theta)|\psi_{0}\rangle$. Two different initial states, $|\psi_{0}\rangle$, are considered, the condensate ansatz (blue, peaked on the left), $|\psi_{c}\rangle=(d_{1}^{\dagger})^{N}|{\rm vac}\rangle/\sqrt{N!}$, and the hard-core ansatz (red, peaked on the right), $|\psi_{h}\rangle=\prod_{j=1}^{N}d_{j}^{\dagger}|{\rm vac}\rangle$. Solid lines show $N=4$ particles in a chain of length $L=8$, while dashed lines have $N=8$ and $L=16$. The pulse durations are selected so that $\theta=\varphi=\pi/2$.
@@
We quantify the accuracy of the gate by the fidelity $|\langle\psi_{X}|\psi\rangle|^{2}$, where $|\psi_{X}\rangle$ is our target state. It is the lowest eigenstate of $H_{X}$. The state produced by the pulse sequence is $|\psi\rangle=U_{Z}U_{X}|\psi_{0}\rangle$. Figure 2 shows the fidelity as a function of $\eta/g_{\rm max}$, using $\delta_{\rm max}=g_{\rm max}$ and $\theta=\varphi=\pi/2$. Exploration of other phase angles can be found in Appendix C. Typical experiments have $g/(2\pi)\sim 10$MHz, yielding a gate time of $T\sim 0.1\mu$s.
@@
To experimentally verify that one has produced the desired state, we propose looking at the time evolution of the density profile under $H_{X}$. Figure 3 shows the required pulse sequence, starting from the product state $|\psi_{0}\rangle$. In this schematic, the $U_{X}$ operation is labeled $\sqrt{\Phi SWAP}$, the $U_{Z}$ operation is labeled “phase gate”, and “probe” labels time evolution under $H_{X}$. (See Sec. V for more discussion of this notation.) The vertical axis shows the magnitude of $g_{\rm c}(t)$ and $\delta_{L}(t)$. The first two pulses correspond to our gate sequence, while the third is used to verify that we have created the desired state. In particular, during this probe pulse the density profile will be static if and only if the system is an eigenstate of $H_{X}$.
To quantify the density evolution during the probe pulse, we introduce the center-of-mass operator $\zeta=(8/L^{2})Z$. If the system is prepared using an imperfect set of gates, e.g. the angle of the phase gate $\varphi$ is not exactly $\pi/2$, the “magnetization” vector $\langle{\bm{M}}\rangle=(\langle X\rangle,\langle Y\rangle,\langle Z\rangle)$ will undergo a Larmor precession around $x$-axis. In the hard-core limit one expects that $\langle\zeta\rangle=\cos(\varphi)\sin(\alpha_{X}t_{h})$, where $\alpha_{X}$ is the coefficient multiplying the hopping term in Eq. (10), and $t_{h}$ is the duration of the probe pulse. Similar results will be found outside of the hard core limit: If one is not in an eigenstate of $H_{X}$, the measured value of $\langle z\rangle$ will reveal pronounced Rabi flops. This structure is shown in Fig. 4 for experimentally relevant parameters. The absence of the Rabi oscillations in the dependence of $\langle\zeta\rangle$ on $t_{h}$ will be indicative of the high quality state preparation.
@@
Figure 4: Expectation value of the center-of-mass operator $\zeta$ obtained numerically for the Bose Hubbard Hamiltonian of a linear chain with $L=8$ and $N=4$, as a function of the duration of the probe pulse in Fig. 3. Here $\eta/2\pi=-270$ MHz, $g_{\rm max}/2\pi=15$MHz, $\delta_{L}/2\pi=25$MHz . The numerical results are in a good agreement with expected sinusoidal dependence predicted in the hardcore limit. Similarly to this prediction, we also see an almost complete suppression of the Rabi oscillations indicative of the high fidelity state preparation. As further explored in Appendix C, this suppression occurs at the optimal $\varphi^{*}=-0.47\pi$ rather than at the nominal $\varphi=-\pi/2$ expected in the hardcore limit. This optimal angle depends on the ratio $\eta/g_{\rm max}$.
@@ III.2 Many-Body Physics of our 2D Model
In the hard core limit $\eta/g\to\infty$, starting with $|\psi_{h}\rangle$, the fidelity approaches 0.51,0.39,0.34 for $L=6,10,15$. This is not as high as the 1D case, as there is no exact mapping of 2D hard core bosons onto non-interacting fermions. Although we do not show it here, the fidelities further drop if one increases $N$. Thus this two dimensional protocol is best reserved for platforms which realize the weakly interacting system.
@@ IV Multi-particle bound states
In the strongly interacting limit $|\eta|\gg|g|,|\delta|$ we can also consider multi-particle bound states where all of the particles sit on the same site – but that bound complex can be on a superposition of sites. These cat states are highly excited eigenstates of the repulsive Bose Hubbard model, and form the true ground state of the attractive model [34, 35]. In what follows we will consider the latter, i.e. use Hamiltonian (1) with $\eta<0$ and $g_{ij}<0$.
We define $|j\rangle$ to be the state where all particles sit on site $j$. Integrating out the configurations where fewer than $N$ particle are on a single site, we can find an effective model Hamiltonian describing dynamics within the manifold of the cat states:
@@
with $g_{ij}$ and $\delta_{i}$ corresponding to the Hamiltonian terms in Eq. (1). The correction terms in $V_{i}$ come from virtual processes where the complex breaks up, and then reforms on its original site.
@@ V Other applications
Variants of the gates discussed here have utility beyond state preparation. In particular, the literature contains a number of works that study $U_{X}(\varphi=\pi)$, for which the phase angle is double the one used for preparing the ground state of the Bose Hubbard model. This $\pi$-gate can be considered a many-particle generalization of the two-qubit iSWAP gate [36]. Given any bit sequence $|\sigma_{1}\sigma_{2}\cdots\sigma_{L}\rangle$, the $U_{X}(\pi)$ gate swaps the order, and applies a phase shift,
@@
Figure 8 shows that for typical parameters the fidelity of the first SWAP is quite large. For $N>1$, subsequent oscillations show smaller amplitudes. This is particularly well illustrated by panel (a), which focuses on the first two swaps. The loss of fidelity is more pronounced for larger particle numbers or weaker interactions.
@@ VI Summary and Outlook
The key concepts used in this work was the mapping between hopping Hamiltonians and either SU(2) or SU(3) spin operators. The SU(2) case is familiar from ‘dual rail encoding’ of quantum spins – where a spin $S/2$ is encoded in the motion of $12S+$ 1 bosons on two sites [50]. The SU(3) case is the generalization to three sites. One could clearly extend this to higher dimensions. Previous theoretical and experimental works have explored engineering hopping matrix elements to implement spin physics or to manipulate quantum information [39, 11, 40, 41, 42, 43, 10, 38, 37, 51].
@@ Appendix C Optimizing the phase gate
In Fig. 9 we show the gate fidelity as a function of $\varphi$, fixing $\theta=\pi/2$. When $\eta/g\to\infty$ the optimal phase angle is $\varphi^{*}=\pi/2$. As one decreases $\eta/gg$ the optimal angle shifts: Fig. 9 shows that $\varphi^{*}$ falls as one moves away from the hard core limit. We understand this result by noting that during time evolution one encounters a spatially dependent density profile (see Fig. (1). For early times the density is higher on the left part of the system, and interactions produce a phase gradient, which must be corrected by adjusting $\varphi$. We find almost no advantage to shifting $\theta$ away from $\pi/2$. This approach can be viewed as a highly specialized variational quantum eigensolver [52].
@@ Appendix D Relationship between eigenstates of $H_{X}$ and eigenstates of uniform hopping Hamiltonian
The central result of our paper is that we can use a multi-site gate to produce the ground state of $H_{X}$ (or $H_{W}$ in 2D). This Hamiltonian is a Bose-Hubbard model with inhomogeneous hopping matrix elements. An obvious concern is that this goal could be perceived as somewhat unnatural. Surely it would be preferable to produce the ground state of Eq. (1) with uniform hoppings. Those concerns are alleviated by noting that, in the strongly interacting limit, the many-body ground state of Eq. (10), $|\psi_{X}\rangle$, has an extremely large overlap with the ground state of the uniform hopping model $|\psi_{\rm u}\rangle$. In Fig. 10 we show the overlap $|\langle\psi_{X}|\psi_{\rm u}\rangle|^{2}$ for $N=L/2$ hard core particles in a chain of length $L$. The overlap is large, even for thousands of sites. One can further aleviate any concerns by following up our gate sequence with an adiabatic evolution step, where the couplings $g_{ij}$ are slowly tuned to a uniform value. We emphasize that even though the $|\psi_{u}\rangle$ is almost identical to $|\psi_{X}\rangle$, the non-uniform hopping is an essential part of the gate. One cannot simply replace $g_{ij}$ with a constant.

### v2

# Multi-site gates for state preparation in quantum simulation of the Bose-Hubbard Model

@@ Abstract
State preparation is one of the most important challenges in analog quantum simulation. The most common approach, adiabatic state preparation, performs poorly in gapless systems and cannot efficiently cross phase boundaries [1, 2, 3]. Driven-dissipative approaches are promising, but require specialized hardware [4, 5]. Here we present a coherent approach to producing the ground state of the Bose-Hubbard model, based upon simple multi-site gates. These are more flexible than the two-site gates used in digital quantum simulation, and allow us to produce the exact ground state, without any Trotter error [6, 7, 8, 9]. Related ideas have been proposed for producing states with applications to quantum information theory [10, 11, 12]. Although our approach can be used in a broad range of architectures [6] it is particularly appealing for simulation based on transmon arrays, as all of the necessary resources are found in the current generation of NISQ era quantum computers[13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Here $d_{i},d_{i}^{\dagger}$ are ladder operators which change the number of excitations present in the transmon on site $i$, $g_{ij}$ parameterizes the strength of the coupling between sites $i$ and $j$, $\delta_{i}$ is the detuning, and $\eta$ encodes the transmon nonlinearity. In modern hardware both $g_{ij}$ and $\delta_{i}$ are tunable, and can be made functions of time [21]. If we interpret $d_{i}$ as the annihilation operator for a boson, Eq. (1) is the Bose-Hubbard model. In that interpretation, $g_{ij}$ is a hopping matrix element, $\delta_{i}$ an on-site potential, and $\eta$ an on-site interaction – more commonly written as $t_{ij},V_{i},$ and $U$. The Bose-Hubbard model is one of the iconic models of strongly interacting quantum matter, displaying fascinating emergent properties, including a phase transition between a superfluid and insulating state[24]. While there are efficient classical methods to calculate any static properties of this model [25], its dynamics are hard to calculate on a classical computer. Thus this is a natural setting to demonstrate quantum advantage [26].
@@ II.1 Single Particle Physics
We begin by presenting a one dimensional model, where $g_{ij}$ vanishes unless $j=i+1$. We will choose these matrix elements so that the hopping part of the Hamiltonian can be mapped onto a SU(2) spin matrix. The structure of spin matrices allow us to construct a gate which produces the superfluid ground state of the Bose-Hubbard model.
@@
In particular, we could start with a particle on the left-most site of the lattice, which is the eigenstate of $Z$ with minimal eigenvalue. It will be mapped onto to the eigenstate of $X$ with minimal eigenvalue.
@@ II.2 Many-Body Physics of our 1D model
Figure 2: Fidelity $|\langle\psi_{X}|\psi\rangle|^{2}$ of state preparation using the pulse sequence in Sec. II.2 for the 1D Bose-Hubbard model. Our target state, $\psi_{X}$, is the ground state of the Bose-Hubbard Hamiltonian $H_{X}$ in Eq. (10, while $|\psi\rangle=U_{Z}(\varphi)U_{X}(\theta)|\psi_{0}\rangle$. Two different initial states, $|\psi_{0}\rangle$, are considered, the condensate ansatz (blue, peaked on the left), $|\psi_{c}\rangle=(d_{1}^{\dagger})^{N}|{\rm vac}\rangle/\sqrt{N!}$, and the hard-core ansatz (red, peaked on the right), $|\psi_{h}\rangle=\prod_{j=1}^{N}d_{j}^{\dagger}|{\rm vac}\rangle$. Solid lines show $N=4$ particles in a chain of length $L=8$, while dashed lines have $N=8$ and $L=16$. The pulse durations are selected so that $\theta=\varphi=\pi/2$.
@@
We quantify the accuracy of the gate by the fidelity $|\langle\psi_{X}|\psi\rangle|^{2}$, where $|\psi_{X}\rangle$ is our target state. It is the lowest eigenstate of $H_{X}$. The state produced by the pulse sequence is
$|\psi\rangle=U_{Z}U_{X}|\psi_{0}\rangle.$ (17)
Figure 2 shows the fidelity as a function of $\eta/g_{\rm max}$, using $\delta_{\rm max}=g_{\rm max}$ and $\theta=\varphi=\pi/2$. Exploration of other phase angles can be found in Appendix C. Typical experiments have $g/(2\pi)\sim 10$MHz, yielding a gate time of $T\sim 0.1\mu$s.
@@
An important feature in Fig. 2 is the size dependence of the fidelity. This can be quantified by finding the value of $|\eta/g_{\rm max}|$ where $|\langle\psi_{X}|\psi\rangle|^{2}=0.5$. We find that this cross-over scale appears to scale as $|\eta/g_{\rm max}|\sim L$ for the hard-core initial conditions and $|\eta/g_{\rm max}|\sim L^{-2}$ for the condensate initial conditions.
@@
To experimentally verify that one has produced the desired state, we propose looking at the time evolution of the density profile under $H_{X}$. Figure 3 shows the required pulse sequence, starting from the product state $|\psi_{0}\rangle$. In this schematic, the $U_{X}$ operation is labeled $\sqrt{\Phi SWAP}$, the $U_{Z}$ operation is labeled “phase gate”, and “probe” labels time evolution under $H_{X}$. (See Sec. V for more discussion of this notation.) The vertical axis shows the magnitude of $g_{\rm c}(t)$ and $\delta_{L}(t)$. The first two pulses correspond to our gate sequence, while the third is used to verify that we have created the desired state. In particular, during this probe pulse the density profile will be static if and only if the system is in an eigenstate of $H_{X}$.
To quantify the density evolution during the probe pulse, we introduce the center-of-mass operator $\zeta=(8/L^{2})Z$. If the system is prepared using an imperfect set of gates, e.g. the angle of the phase gate $\varphi$ is not exactly $\pi/2$, the “magnetization” vector $\langle{\bm{M}}\rangle=(\langle X\rangle,\langle Y\rangle,\langle Z\rangle)$ will undergo a Larmor precession around $x$-axis. In the hard-core limit one expects that $\langle\zeta\rangle=\cos(\varphi)\sin(\alpha_{X}t_{h})$, where $\alpha_{X}$ is the coefficient multiplying the hopping term in Eq. (10), and $t_{h}$ is the duration of the probe pulse. Similar results will be found outside of the hard core limit: If one is not in an eigenstate of $H_{X}$, the measured value of $\langle z\rangle$ will reveal pronounced Rabi flops. This structure is shown in Fig. 4 for experimentally relevant parameters. The absence of the Rabi oscillations in the dependence of $\langle\zeta\rangle$ on $t_{h}$ will be indicative of the high quality state preparation. For comparison, the fidelity is a quadratic function of $\varphi$, peaked at $\varphi=0$.
@@
Figure 4: Expectation value of the center-of-mass operator $\zeta$ obtained numerically for the Bose-Hubbard Hamiltonian of a linear chain with $L=8$ and $N=4$, as a function of the duration of the probe pulse in Fig. 3. Here $\eta/2\pi=-270$ MHz, $g_{\rm max}/2\pi=15$MHz, $\delta_{L}/2\pi=25$MHz . The numerical results are in a good agreement with expected sinusoidal dependence predicted in the hardcore limit. Similarly to this prediction, we also see an almost complete suppression of the Rabi oscillations indicative of the high fidelity state preparation. As further explored in Appendix C, this suppression occurs at the optimal $\varphi^{*}=-0.47\pi$ rather than at the nominal $\varphi=-\pi/2$ expected in the hardcore limit. This optimal angle depends on the ratio $\eta/g_{\rm max}$.
@@ III.2 Many-Body Physics of our 2D Model
In the hard core limit $\eta/g\to\infty$, starting with $|\psi_{h}\rangle$, the fidelity approaches 0.51,0.39,0.34 for $L=6,10,15$. This is not as high as the 1D case, as there is no exact mapping of 2D hard core bosons onto non-interacting fermions. For fixed $L$ the fidelity is a non-monotonic function of $N$. In the hard-core limit it is largest when either the density of particles or holes is small. We would argue that this two dimensional protocol is best reserved for platforms which realize the weakly interacting system.
@@ IV Multi-particle bound states
In the strongly interacting limit $|\eta|\gg|g|,|\delta|$ we can also consider multi-particle bound states where all of the particles sit on the same site – but that bound complex can be on a superposition of sites. These cat states are highly excited eigenstates of the repulsive Bose-Hubbard model, and form the true ground state of the attractive model [34, 35]. In what follows we will consider the latter, i.e. use Hamiltonian (1) with $\eta<0$ and $g_{ij}<0$.
We define $|j\rangle$ to be the state where all particles sit on site $j$. Integrating out the configurations where fewer than $N$ particle are on a single site, we use $N$’th order perturbation theory to find an effective model Hamiltonian describing dynamics within the manifold of the cat states:
@@
with $g_{ij}$ and $\delta_{i}$ corresponding to the Hamiltonian terms in Eq. (1). The correction terms in $V_{i}$ come from virtual processes where the complex breaks up, and then reforms on its original site. See [34] for a full derivation.
@@ V Other applications
Variants of the gates discussed here have utility beyond state preparation. In particular, the literature contains a number of works that study $U_{X}(\varphi=\pi)$, for which the phase angle is double the one used for preparing the ground state of the Bose-Hubbard model. This $\pi$-gate can be considered a many-particle generalization of the two-qubit iSWAP gate [36]. Given any bit sequence $|\sigma_{1}\sigma_{2}\cdots\sigma_{L}\rangle$, the $U_{X}(\pi)$ gate swaps the order, and applies a phase shift,
@@ VI Summary and Outlook
The key concepts used in this work was the mapping between hopping Hamiltonians and either SU(2) or SU(3) spin operators. The SU(2) case is familiar from ‘dual rail encoding’ of quantum spins – where a spin $S/2$ is encoded in the motion of $12S+$ 1 bosons on two sites [50]. The SU(3) case is the generalization to three sites, resulting in a 2D triangular lattice with coordination number 6. One could clearly extend this to higher dimensions. Using $SU(4)$ we can produce a model on a 3-dimensions cubic closed pack (face centered cubic) lattice with coordination number 12, forming a tetrahedron, similar to a stack of oranges in a grocery store. The sites are labeled by integers $(n_{a},n_{b},n_{c},n_{d})$ with $n_{a}+n_{b}+n_{c}+n_{d}=\ell$. There are no obvious generalizations to lattices with other coordination numbers. Previous theoretical and experimental works have explored engineering hopping matrix elements to implement spin physics or to manipulate quantum information [39, 11, 40, 41, 42, 43, 10, 38, 37, 51].
@@ Appendix C Optimizing the phase gate
In Fig. 9 we show the gate fidelity as a function of $\varphi$, fixing $\theta=\pi/2$. When $\eta/g\to\infty$ the optimal phase angle is $\varphi^{*}=\pi/2$. As one decreases $\eta/gg$ the optimal angle shifts: Fig. 9 shows that $\varphi^{*}$ falls as one moves away from the hard core limit. We understand this result by noting that during time evolution one encounters a spatially dependent density profile (see Fig. (1). For early times the density is higher on the left part of the system, and interactions produce a phase gradient, which must be corrected by adjusting $\varphi$. We find almost no advantage to shifting $\theta$ away from $\pi/2$. This approach can be viewed as a highly specialized variational quantum eigensolver [52], or an application of the Quantum Approximate Optimization Algorithm [53].
@@ Appendix D Relationship between eigenstates of $H_{X}$ and eigenstates of uniform hopping Hamiltonian
The central result of our paper is that we can use a multi-site gate to produce the ground state of $H_{X}$ (or $H_{W}$ in 2D). This Hamiltonian is a Bose-Hubbard model with inhomogeneous hopping matrix elements. An obvious concern is that this goal could be perceived as somewhat unnatural. Surely it would be preferable to produce the ground state of Eq. (1) with uniform hoppings. Those concerns are alleviated by noting that, in the strongly interacting limit, the many-body ground state of Eq. (10), $|\psi_{X}\rangle$, has an extremely large overlap with the ground state of the uniform hopping model $|\psi_{\rm u}\rangle$. In Fig. 10 we show the overlap $|\langle\psi_{X}|\psi_{\rm u}\rangle|^{2}$ for $N=L/2$ hard core particles in a chain of length $L$. The overlap is large, even for thousands of sites. One can further alleviate any concerns by following up our gate sequence with an adiabatic evolution step, where the couplings $g_{ij}$ are slowly tuned to a uniform value. The conditions for adiabaticity are less stringent here, compared with transforming an insulator into a superfluid, as the initial and final state are closer together.
We emphasize that even though the $|\psi_{u}\rangle$ is almost identical to $|\psi_{X}\rangle$, the non-uniform hopping is an essential part of the gate. One cannot simply replace $g_{ij}$ with a constant.